# Towards Adaptive Visual Token Pruning for Large Multimodal Models

### v1

@@ Introduction
Image Language Understanding
We conduct a comprehensive evaluation of our visual token pruning approach against several baselines. As shown in Table 1, our method consistently outperforms prior works across a range of token retention ratios on LLaVA-1.5-7B. At a moderate retention of 192 tokens, our approach achieves an MME perception score of 1467.30, with only a 2.7% drop relative to the dense model, compared to 13.7% and 16.2% drops for FastV and ToMe, respectively. Compared with the strongest baseline, DivPrune, our method improves MMB accuracy from 62.28 to 63.75, a relative gain of 2.4%. Moreover, the standard deviation of our results is smaller than that of other methods; for example, on ChartQA, our method achieves a standard deviation of 0.75, compared to 0.78 for DivPrune. To further assess the scalability of our approach on larger models, we extend our evaluation to LLaVA-1.5-13B and conduct a comparative analysis against DivPrune under varying visual token retention settings. As presented in Table 2, our method consistently surpasses DivPrune, demonstrating its effectiveness at higher model scale. In particular, on the MMB benchmark, our method achieves a notable score of 67.35 with only 144 retained tokens, outperforming DivPrune’s 66.41 under the same setting. These results highlight the generalization ability of our pruning strategy in large scale settings.
In addition, we conduct experiments on LLaVA-NEXT-7B using a larger pool of visual tokens across different retention settings. As shown in Table 3, our method consistently achieves the strongest performance across a range of compression levels. Furthermore, our experiments on Qwen2-VL-7B also demonstrate the effectiveness of the proposed method (see Appendix 5).
@@
Video Language Understanding
In this section, we evaluate our method on the video based multimodal model LLaVA-NEXT-Video-7B. Using the same pruning strategy as in our image understanding experiments, we conduct evaluations on the COCO dataset and report standard image captioning metrics, including BLEU-1/2/3/4, ROUGE-L and CIDEr. We compare our method against DivPrune, a strong baseline for visual token pruning. As shown in Table 4, our approach consistently outperforms DivPrune across all evaluation metrics under various token retention settings. For instance, with 960 tokens retained, our method improves the BLEU-4 score from 28.46 to 30.41, representing a relative gain of 6.9%. These results demonstrate the robustness of our method in video scenarios.
Efficiency Analysis (Accounting for Pruning Overhead)
In this section, we evaluate the inference efficiency of our method on LLaVA-NEXT-7B using the MME dataset. We repeat the process three times and report the average result. As shown in Table 5, we report the average inference time (time to generate the complete output) per sample with a batch size of 1 and 288 visual tokens retained. Compared to the original dense model (15.53 GB, 235.2 ms), all pruning methods significantly reduce memory usage and cut inference latency by more than half. Our method achieves an inference time of 100.13 ms, representing a 57% reduction compared to the dense baseline. When compared with other pruning baselines, our approach achieves lower latency than VTW and FastV. Although our latency is slightly higher than DivPrune, this small gap is negligible given the superior task performance consistently achieved under the same token budget. It is worth noting that pruning is performed during inference; therefore, the reported inference latency includes the overhead introduced by the pruning process. These results demonstrate that our pruning strategy provides substantial improvements in memory and latency efficiency.
Method Memory (G) Latency (ms)
Dense 15.53 235.2
VTW 13.63 103.71
FastV 13.63 108.29
DivPrune 13.63 99.53
Ours 13.63 100.13
Table 5: Comparison of peak GPU memory usage and average per-sample latency for our pruning method and baselines on LLaVA-NEXT-7B, evaluated on the MME dataset with a batch size of 1 and 288 retained visual tokens.
Ablation Study
Cross-Modal Token Filtering.
To assess the effectiveness of the proposed cross-modal alignment aware token filtering strategy, we perform an ablation study on LLaVA-1.5-7B across several representative benchmarks. As shown in Figure 2, we evaluate three configurations: the original dense model, a pruning variant without cross-modal filtering, and our pruning approach incorporating cross-modal filtering. The experimental results clearly demonstrate that our method successfully eliminates redundant visual tokens while minimizing performance degradation. Compared to the dense baseline, our approach consistently maintains competitive performance across all benchmarks, exhibiting only marginal declines. In contrast, the pruning strategy without cross-modal filtering suffers from noticeably larger performance drops, underscoring the critical role of cross-modal alignment in preserving model performance during token compression.
Figure 2: Ablation study on the impact of cross-modal alignment aware token filtering in LLaVA-1.5-7B across multiple benchmarks.
Additionally, we conduct ablation studies to validate the effectiveness of our intra-modal representation maximization strategy and the $L_{2}$ norm based mutual information approximation (Appendix 2).
Additional Results
We conduct a Hyperparameter Analysis, such as varying $N_{1}$, and observe consistently strong performance across different settings, demonstrating the robustness of our method (Appendix 3). Moreover, we perform a Case Study to show that the pruned model retains its ability to generate high quality content (Appendix 4).
Conclusion
[…]

### v2

@@ 5.1 Experimental Setup
5.2 Main Results
We conduct a comprehensive evaluation of our visual token pruning approach against several baselines. As shown in Table 1, our method consistently outperforms prior works across a range of token retention ratios on LLaVA-1.5-7B. At a moderate retention of 192 tokens, our approach achieves an MME perception score of 1467.30, with only a 2.7% drop relative to the dense model, compared to 13.7% and 16.2% drops for FastV and ToMe, respectively. Compared with the strongest baseline, DivPrune, our method improves MMB accuracy from 62.28 to 63.75, a relative gain of 2.4%. To further assess the scalability of our approach on larger models, we extend our evaluation to LLaVA-1.5-13B and conduct a comparative analysis against VisionZip and DivPrune under varying visual token retention settings. As presented in Table 2, our method consistently surpasses others, demonstrating its effectiveness at higher model scale. These results highlight the generalization ability of our pruning strategy in large scale settings.
In addition, we conduct experiments on LLaVA-NEXT-7B using a larger pool of visual tokens across different retention settings. As shown in Table 3, our method consistently achieves the strongest performance across a range of compression levels. Furthermore, our experiments on LLaVA-NEXT-Video-7B and Qwen2-VL-7B also demonstrate the effectiveness of the proposed method (see Appendix C).
@@ 5.2 Main Results
5.3 Efficiency Analysis (Accounting for Pruning Overhead)
In this section, we evaluate the inference efficiency of our method on LLaVA-NEXT-7B using the MME dataset. We repeat the process three times and report the average result. For a fair comparison, the number of decoding steps for each method is fixed to the minimum decoding steps required among all methods. As shown in Table 4, we report the average inference time (time to generate the complete output) per sample with a batch size of 1 and 288 visual tokens retained. Compared to the original dense model (15.53 GB, 235.2 ms), all pruning methods significantly reduce memory usage and cut inference latency by more than half. Our method achieves an inference time of 100.13 ms, representing a 57% reduction compared to the dense baseline. When compared with other pruning baselines, our approach achieves lower latency than VTW and FastV. Although our latency is slightly higher than DivPrune, this small gap is negligible given the superior task performance consistently achieved under the same token budget. It is worth noting that pruning is performed during inference; therefore, the reported inference latency includes the overhead introduced by the pruning process. These results demonstrate that our pruning strategy provides substantial improvements in memory and latency efficiency.
Method Memory (G) Latency (ms)
Dense 15.53 235.2
VTW 13.63 103.71
FastV 13.63 108.29
DivPrune 13.63 99.53
Ours 13.63 100.13
Table 4: Comparison of peak GPU memory usage and average per-sample latency for our pruning method and baselines on LLaVA-NEXT-7B, evaluated on the MME dataset with a batch size of 1 and 288 retained tokens.
5.4 Comparison with our DivPrune variants (DivPrune)
We further extend the DivPrune approach to enable a more comprehensive comparison. Concretely, we integrate our cross-modal pruning strategy with DivPrune under its default configuration, forming a new variant denoted as DivPrune*. We evaluate this variant on LLaVA-1.5-7B while retaining 64 visual tokens. As shown in Table 5, our method still surpasses DivPrune*. Combined with the earlier results, we also observe that DivPrune* achieves noticeable gains over the original DivPrune, highlighting the crucial role of cross-modal alignment in boosting performance. Moreover, these findings demonstrate that our greedy intra-modal expectation representation maximization strategy is more effective than the minimum distance maximization employed in DivPrune.
Method ChartQA COCO MMB MME MMU NoCaps
DivPrune 15.98 1.01 60.28 1355.28 36.02 0.96
Ours 16.36 1.02 61.34 1359.12 36.11 0.98
Table 5: Comparison of DivPrune variants (DivPrune) with our method on LLaVA-1.5-7B across multiple benchmarks.
5.5 Ablation Study
To further validate the effectiveness of our approach, we conduct ablation studies with three different configurations. $ablation_{1}$ swaps the order of the two pruning stages by applying intra-modal representation maximization before cross-modal alignment. $ablation_{2}$ removes intra-modal pruning and only retains cross-modal alignment based pruning. $ablation_{3}$ eliminates cross-modal alignment pruning and applies only intra-modal representation maximization. As shown in Table 6, the performance consistently drops once any component is removed, indicating the necessity of each part of our framework. Moreover, using only cross-modal pruning or altering the pruning order leads to more severe degradation. We conjecture that this is because text-related visual information is often concentrated in local regions, where tokens align with the text but may carry highly similar semantics. Additional experiments in Section E further demonstrate that with properly chosen pruning parameters for each stage, our method achieves strong performance.
Method ChartQA COCO MMB MME MMU NoCaps
$ablation_{1}$ 13.12 0.83 50.00 1114.77 34.56 0.76
$ablation_{2}$ 13.08 0.83 49.57 1108.64 34.33 0.75
$ablation_{3}$ 16.16 1.00 59.97 1311.46 36.02 0.97
Ours 16.36 1.02 61.34 1359.12 36.11 0.98
Table 6: Ablation study of different pruning strategies on LLaVA-1.5-7B across multiple benchmarks.
5.6 Additional Results
We conduct a Hyperparameter Analysis, such as varying $N_{1}$, and observe consistently strong performance across different settings, demonstrating the robustness of our method (Appendix E). Moreover, we perform a Case Study to show that the pruned model retains its ability to generate high quality content (Appendix F).
6 Conclusion
Limitations
In this work, we conduct extensive experiments to evaluate the effectiveness of our multimodal model token pruning method. The results demonstrate that our approach achieves competitive performance compared to the baselines. However, due to computational constraints, we have not yet been able to evaluate it on larger scale models, such as those with 70 billion parameters. Exploring the scalability of our method to such large models constitutes an important direction for future work.
Appendix A Token Redundancy Across Modalities
Previous studies Shang et al. (2024a); Alvar et al. (2025); Li et al. (2025a) primarily decide to prune visual tokens based on the discrepancy in the number of visual and textual tokens. However, these works do not thoroughly analyze the differences in redundancy across modalities. In this work, we provide a more detailed pruning guidance by examining the redundancy of visual and textual tokens from three perspectives: attention mechanisms, semantic distribution and information repetition. Our findings reveal that visual tokens exhibit substantially higher redundancy than textual tokens. Motivated by this observation, we retain all textual tokens during the pruning process and apply pruning exclusively to visual tokens.
The attention distribution in multimodal models is highly skewed, with textual tokens receiving significantly higher weights than visual tokens Li et al. (2025b); Lee et al. (2025). This suggests that the model primarily relies on text for semantic understanding. Visual inputs often consist of hundreds or thousands of tokens, far exceeding the number of textual tokens. Given the quadratic complexity of attention with respect to token count, this mismatch between quantity and contribution is pronounced. Despite the large number of visual tokens, their impact on the final representation is limited, resulting in poor trade offs between computational cost and semantic gain and reflecting the high redundancy of visual tokens.
Each token in text typically carries explicit semantic meaning, such as a noun, verb, or conjunction, reflecting natural language as a highly optimized discrete encoding system with dense and relatively uniform information distribution Hirschberg and Manning (2015). In contrast, semantic information in images is highly concentrated in a small number of salient regions, such as foreground objects, while the majority of visual tokens correspond to low-semantic areas like sky or walls, which mostly contain low-frequency textures or superficial variations and contribute little to high-level semantic understanding. Consequently, the effective semantic density of visual tokens is significantly lower than that of textual tokens.
Compared to textual tokens, visual tokens exhibit higher information redundancy due to differences in their generation processes and underlying information structures. Textual tokens, derived from natural language, follow Zipf’s law Newman (2005); Adamic and Huberman (2002) and represent distinct semantic units with low repetition. In contrast, visual tokens are typically generated by uniformly partitioning images or extracting low-level features without semantic organization. For example, a sky region is divided into multiple similar patches that cluster in embedding space and jointly represent a single high-level concept, causing redundancy. Visual tokenization resembles physical partitioning rather than frequency-driven semantic abstraction, forcing models to process many similar tokens and resulting in computational and representational inefficiency.
Appendix B Proof of $L_{2}$ Norm Approximation to Mutual Information
When considering two continuous random variables $x_{i}^{v}$ and $x_{j}^{t}$, corresponding to two distributions, the mutual information between them is defined as follows:
$I(x_{i}^{v};x_{j}^{t})=\int p(x_{i}^{v},x_{j}^{t})\log\frac{p(x_{i}^{v}\mid x_{j}^{t})}{p(x_{i}^{v})}\,dx_{i}^{v}dx_{j}^{t}$
To render this quantity tractable, we make the following assumptions:
  • The conditional distribution $p(x_{i}^{v}\mid x_{j}^{t})$ is modeled as an isotropic Gaussian. We conduct extensive experiments and find that, during the fitting process, the conditional distribution exhibits per-dimension means and standard deviations concentrated around (-0.0035) and (0.7842), respectively, which substantiates the validity of our hypothesis. Then, we can derive the following expression:
    $p(x_{i}^{v}\mid x_{j}^{t})=\mathcal{N}(x_{j}^{t},\sigma^{2}I)$ (18)
    Here, $I$ denotes the identity matrix.
  • The marginal distribution $p(x_{i}^{v})$ is generally unknown and intractable; thus, we approximate it as a constant $C_{2}$, following common practice in variational inference and contrastive learning.
Based on the above assumptions, we obtain the following expression:
$\log p(x_{i}^{v}\mid x_{j}^{t})=C_{1}-\frac{1}{2\sigma^{2}}\|x_{i}^{v}-x_{j}^{t}\|_{2}^{2}$ (19)
$I(x_{i}^{v};x_{j}^{t})\approx-\frac{1}{2\sigma^{2}}\,\mathbb{E}_{p(x_{i}^{v},x_{j}^{t})}\!\left[\|x_{i}^{v}-x_{j}^{t}\|_{2}^{2}\right]+C$
Here, $C_{1}=-\frac{d}{2}\log(2\pi\sigma^{2})$ is a constant independent of the inputs, and we denote $C=C_{1}-C_{2}$. Therefore, the $L_{2}$ norm provides a theoretically grounded approximation of mutual information.
Appendix C Additional Comparative Experiments
We evaluate our method on the video based multimodal model LLaVA-NEXT-Video-7B. Using the same pruning strategy as in our image understanding experiments, we conduct evaluations on the COCO dataset and report standard image captioning metrics, including BLEU-1/2/3/4, ROUGE-L and CIDEr. We compare our method against DivPrune, a really strong baseline for visual token pruning. As shown in Table 7, our approach consistently outperforms DivPrune across all evaluation metrics under various token retention settings. For instance, with 960 tokens retained, our method improves the BLEU-4 score from 28.46 to 30.41, representing a relative gain of 6.9%. These results demonstrate the robustness of our method in video scenarios.
@@ Appendix C Additional Comparative Experiments
To more comprehensively assess the effectiveness of our method, we further perform pruning experiments on Qwen2-VL-7B, pruning 88.9% of tokens. We test our approach against VisionZip and DivPrune across multiple datasets. As illustrated in Table 8, our method achieves superior performance, underscoring its robustness and effectiveness even under extreme pruning scenarios.
Method MMB MME VQATEXT
Dense 78.26 1894 65
VisionZip 74.26 1702.97 57.18
DivPrune 73.96 1699.03 58.29
Ours 74.90 1706.74 58.43
Table 8: Performance comparison of 88.9% token pruning on Qwen2-VL-7B.
Appendix D Metric Analysis
In addition to the theoretical analysis of approximating mutual information (MI) using the $L_{2}$ norm, we conduct empirical validation on LLaVA-1.5-7B with 64 visual tokens retained. Specifically, we implement cross-modal alignment using three metrics: mutual information, its $L_{2}$ norm approximation, and cosine similarity. We use scikit-learn Pedregosa et al. (2011) to compute mutual information, where the library employs a nonparametric $k$ nearest neighbors approximation method, with $k$ set to 3 by default. As shown in Table 9, the method using the $L_{2}$ norm approximation achieves performance comparable to that of exact mutual information, and consistently outperforms the one using cosine similarity. We guess that this phenomenon occurs because semantic alignment between textual and visual tokens requires not only directional consistency but also a substantial overlap in their feature distributions. We further compare performance under different intra-modal redundancy metrics, namely cosine similarity and the $L_{2}$ norm. As shown in Table 10, cosine similarity consistently yields better results. This suggests that cosine similarity is generally more effective for measuring redundancy within unimodal data, which is consistent with prior empirical findings in the literature.
Metric ChartQA COCO MMB MME MMU NoCaps
MI 16.38 1.02 61.38 1362.88 36.12 0.98
$L_{2}$ 16.36 1.02 61.34 1359.12 36.11 0.98
cos 16.26 1.01 61.15 1357.05 35.83 0.97
Table 9: Performance comparison among various cross-modal alignment metrics across multiple benchmarks.
Metric ChartQA COCO MMB MME MMU NoCaps
$L_{2}$ 16.32 1.0 61.26 1356.91 36.02 0.98
cos 16.36 1.02 61.34 1359.12 36.11 0.98
Table 10: Performance comparison among various intra-modal redundancy metrics across multiple benchmarks.
Appendix E Hyperparameter Analysis
In this study, we conduct an analysis of the influence of the token retention parameter $N_{1}$ in the first stage cross-modal token pruning process on the overall performance of our method. As presented in Table 11, we evaluate several representative configurations on LLaVA-1.5-7B, including $N_{1}=\left\lfloor 0.9N\right\rceil$, $N_{1}=\left\lfloor 0.8N\right\rceil$, $N_{1}=\left\lfloor 0.75N\right\rceil$, and $N_{1}=\left\lfloor 0.7N\right\rceil$, where $N$ represents the total number of visual tokens in the dense model. In all cases, the final number of retained tokens is fixed at 64. Our method consistently achieves strong and stable performance across these different settings, indicating that it is largely insensitive to the choice of $N_{1}$. These strong results highlight the robustness of our method to different hyperparameter settings.
$N_{1}$ ChartQA COCO MMB MME MMU NoCaps
$\left\lfloor 0.9N\right\rceil$ 16.48 1.02 60.99 1332.41 36.11 0.97
$\left\lfloor 0.8N\right\rceil$ 16.36 1.02 61.34 1359.12 36.11 0.98
$\left\lfloor 0.75N\right\rceil$ 16.28 1.02 61.08 1334.68 36.11 0.97
$\left\lfloor 0.7N\right\rceil$ 16.00 1.02 61.43 1347.34 35.78 0.97
Table 11: Impact of different values of the parameter $N_{1}$ on model performance using LLaVA-1.5-7B.
Appendix F Case Study
To qualitatively assess the effectiveness of our pruning strategy, we conduct a case study using LLaVA-1.5-7B on samples from the COCO dataset. As illustrated in Figure 2, we compare the captions generated by the original dense model and our pruned model, which retains significantly fewer visual tokens. The results show that the pruned model produces captions that remain semantically coherent and fluent, closely aligning with those generated by the dense model. This demonstrates that our method effectively removes redundant visual tokens while preserving caption generation capabilities. Overall, the case study highlights that substantial token reduction can be achieved without compromising output quality, underscoring the practical utility of our approach.
Figure 2: Case study comparing captions generated by dense and pruned models (LLaVA-1.5-7B) on the COCO dataset, demonstrating output consistency despite substantial token reduction.
Reproducibility statement
We strive to ensure the reproducibility of our results. Full details are provided in the main paper and the appendix. Our implementation is built on PyTorch and standard open-source libraries. We provide key code implementations to facilitate reproducibility and further research.